8. Describe all non-trivial proper subgroups of the direct product Z/27 x Z/2Z of groups. ANSWER:
Answer #1

Soluhon 2/22 = {,73 where Ō = 0+ 22 T = 1 + 22 => 420 x 2/20 ={(510), 1017), (1ā), (1,7)} Thus order of 2x 2/2 is 4 Divisors